yourgrandpa:

noelle spends $6.87 on materials for each bracelet she makes. she sells the bracelets at a craft fair for $18 each. what is the percent markup on the bracelets? A)38.2% B)61.8% C)162.0% D)362.0%

AZ:

\( \text{% markup = }\dfrac{\text{selling price - cost to make it}}{\text{cost to make it}} \times 100\)

yourgrandpa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ \( \text{% markup = }\dfrac{\text{selling price - cost to make it}}{\text{cost to make it}} \times 100\) \(\color{#0cbb34}{\text{End of Quote}}\) 687.00 ?

AZ:

6.87 is the cost to make it you're selling it for 18 so plug it in (18-6.87)/6.87 * 100 = ?

yourgrandpa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ 6.87 is the cost to make it you're selling it for 18 so plug it in (18-6.87)/6.87 * 100 = ? \(\color{#0cbb34}{\text{End of Quote}}\) 162?

AZ:

And that's your percent markup!

yourgrandpa:

thankss(:

AZ:

np!!(: